OpenStudy (anonymous):

transform these equation to rectangular form r= 3 cos theta-2 sin theta

OpenStudy (anonymous):

@ leo302 use the formula sin Q=y/r cos Q=x/r and r^2=x^2 + y^2, now r= 3 cos Q-2 sin Q r=3x/r -2y/r r=(3x-2y)/r r^2=3x-2y=x^2+y^2 or 3x-2y=x^2+y^2 or x^2+y^2 -3x+2y=0 can you do the next step?